OpenStudy (anonymous):

Write in standard form the equation of the line passing through the given point with the given slope. Slope = -8, point= (-2,-2)

OpenStudy (anonymous):

8x+y=14 i believe

OpenStudy (anonymous):

y-(-2)=-8(x--2) =y+2=-8(x+2) =y+2=-8x-16 =y=-8x-18 divide by2 so y=-4x-8